# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

### v1

@@ C.1 Details of the Tango loss
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-\mathbb{E}_{x_{t}}\langle P(x_{t})s_{\theta}(x_{t},t),P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})\rangle+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-\int\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}p_{\sigma_{t}}(x_{t})\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-\mathbb{E}_{x}\int\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\int p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\right\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-\int\int\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-\mathbb{E}_{x,x_{t}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle+C$
$=$ $\mathbb{E}_{x,x_{t}}\|s^{\parallel}_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}+C_{1}.$
Therefore,
$\arg\min_{s_{\theta}^{\parallel}}\mathbb{E}_{x,x_{t}}\|s_{\theta}^{\parallel}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}=P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}).$ (74)
The overall loss calculation is summarized in Algorithm 1.
@@ C.2 Details of the sampling algorithm
0:  trained neural network $s_{\theta}$, total number of discrete SDE steps $N$, projection operator $\pi$
1:  $t\leftarrow T$, $\Delta t>0=T/N$
2:  $x_{t}\sim\mathcal{N}(0,\sigma_{\max}^{2}I)$
@@
0:  Trained neural network $s_{\theta}$, total number of discrete SDE steps $N$, threshold $\tilde{\sigma}$, number of steps $n_{0}$ and step size $\alpha_{\mathrm{ld}}$ for the Langevin dynamics on manifolds, projection operator $\pi$
1:  $t\leftarrow T$, $\Delta t>0=T/N$, $x_{t}\sim\mathcal{N}(0,\sigma_{\max}^{2}I)$
@@
Appendix D Experiment details
The details of each experiment are provided in the subsections below. The neural network used is a multilayer perceptron (MLP) with SiLU as the activation function. Models are trained using PyTorch, utilizing the Adam optimizer with a fixed learning rate and gradients are clipped when their $2$-norm exceeds a predefined threshold. An exponential moving average of the model weights [24] is applied with a decay rate of $0.999$. In each run, the dataset is divided into training and test sets with ratio $8:2$. All experiments are conducted on either a single Tesla V100-PCIE-32GB GPU or an NVIDIA A40 GPU with 48GB of memory. The values of all parameters used in the experiments are listed in Table 5.
@@ Appendix D Experiment details
Table 5: Parameters in our experiments. $\sigma_{\min}$, $\sigma_{\max}$, and $T$ are the parameters of the SDE. $c_{\mathrm{niso}}$ and $c_{\mathrm{tango}}$ are the parameters for Niso-DM and Tango-DM, respectively. $N$, $n_{0}$, and $\alpha_{ld}$ are the hyperparameters of the sampling algorithm. $N_{\mathrm{epoch}}$, $B$, $\mathrm{lr}$, and $\mathrm{clip}$ denote the total epochs, batch size, learning rate, and gradient clipping threshold during training. $N_{\mathrm{node}}$ and $N_{\mathrm{layer}}$ represent the number of hidden nodes per layer and the number of hidden layers in the neural networks, respectively. $\mathcal{D}$ indicates the dataset size.
@@
D.1 The hyperplane in 3D space
@@
The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. Specifically, the means of the nine modes are $(-1,-1)$, $(-1,0)$, $(-1,1)$, $(0,-1)$, $(0,0)$, $(0,1)$, $(1,-1)$, $(1,0)$, and $(1,1)$, and each Gaussian has a standard deviation of $0.3$.
D.2 Mesh data
@@
To create the datasets, we first obtain the clamped eigenfunction of the Laplacian operator on a mesh that has been upsampled threefold for the original mesh. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs. For the distribution on the cow mesh, we additionally discarded the points on the horns and tail.
D.3 High-dimensional special orthogonal group
@@
The manifold $\mathrm{SO}(10)$ is defined as $\{Q\in\mathbb{R}^{10\times 10}|QQ^{T}=I_{10},\operatorname{det}(Q)=1\}$, which represents (a connected component of) the zero-level set of the map $\xi:\mathbb{R}^{100}\rightarrow\mathbb{R}^{55}$. The components of $\xi$ correspond to the upper triangular portion of the matrix $Q^{T}Q-I_{10}$. The dataset is constructed as a mixture of 5 wrapped normal distributions. Each wrapped normal distribution is the image (under the exponential map) of a normal distribution defined in the tangent space at a pre-selected center. For more details, refer to [17].
D.4 Alanine dipeptide
@@
To generate the dataset, we follow [17] to obtain the target distribution on the manifold =$\mathcal{M}=\{x\in\mathbb{R}^{30}|\phi(x)=-70^{\circ}\}$. We also ensure that the generated distribution is $\mathrm{SE}(3)$-invariant (i.e., invariant under rotations and translations) by incorporating alignment before feeding the data into the neural network.
Appendix E Bias-variance tradeoff
@@
For Niso-DM, the singularity is governed by the additional noise scale $c_{\mathrm{niso}}$ for Niso-DM, rather than $\sigma_{\min}$. As a result, our method exhibits greater robustness as $\sigma_{\min}$ decreases. The blue line in Figure 3(a) demonstrates the stability of the error for Niso-DM under varying $\sigma_{\min}$.
Figures 3(b) and 3(c) show the impact of $c_{\mathrm{niso}}$ in Niso-DM and $c_{\mathrm{tango}}$ in Tango-DM on the error. The selection of these hyperparameters also involves a bias-variance trade-off.

### v3

@@ Appendix C Algorithm details
Figure 2 illustrates the schematic diagrams of the vanilla method, along with our proposed Niso-DM and tango-DM.
Figure 2: Schematic diagrams of the vanilla method (left), our proposed Niso-DM (middle), and tango-DM (right).
@@ C.1 Details of the Tango loss
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\mathbb{E}_{x_{t}}\langle P(x_{t})s_{\theta}(x_{t},t),P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t})\rangle+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}p_{\sigma_{t}}(x_{t})\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\int_{\mathcal{M}}p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\right\rangle\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\int_{\mathbb{R}^{n}}\int_{\mathcal{M}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle p_{\sigma_{t}}(x_{t}|x)p_{0}(x)\mathrm{d}x\mathrm{d}x_{t}+C$
$=$ $\mathbb{E}_{x_{t}}\|P(x_{t})s_{\theta}(x_{t},t)\|^{2}-2\mathbb{E}_{x,x_{t}}\left\langle P(x_{t})s_{\theta}(x_{t},t),\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\right\rangle+C$
$=$ $\mathbb{E}_{x,x_{t}}\|s^{\parallel}_{\theta}(x_{t},t)-P(x_{t})\nabla_{x_{t}}\log p_{\sigma_{t}}(x_{t}|x)\|^{2}+C_{1}\,$,
where $C$ and $C_{1}$ are constants independent of $\theta$. Therefore, the minimizer of the loss (20) satisfies that $s_{\theta}^{\parallel}(x,t)=P(x)\nabla_{x}\log p_{\sigma_{t}}(x)$. The overall loss calculation is summarized in Algorithm 1.
@@ C.2 Details of the sampling algorithm
0: trained neural network $s_{\theta}$, total number of discrete SDE steps NN, projection operator $\pi$
[…]
@@
0: trained neural network $s_{\theta}$, total number of discrete SDE steps NN, threshold $\tilde{\sigma}$, number of steps $n_{0}$ and step size $\alpha_{\mathrm{ld}}$ for the Langevin dynamics on manifolds, projection operator $\pi$
[…]
@@
Appendix D Experimental details
The details of each experiment are provided in the subsections below. The neural networks used are multilayer perceptrons (MLPs) with SiLU activations. Models are trained using PyTorch, utilizing the Adam optimizer with a fixed learning rate, and gradients are clipped when their $2$-norm exceeds a predefined threshold. An exponential moving average of the model weights [30] is applied with a decay rate of $0.999$. In each run, the dataset is divided into training and test sets with a ratio $8:2$. All experiments are conducted on either a single Tesla V100-PCIE-32GB GPU or an NVIDIA A40 GPU with 48 GB of memory. The values of all parameters used in the experiments are listed in Table 4.
@@ Appendix D Experimental details
Table 4: Parameters in our experiments. $\sigma_{\min}$, $\sigma_{\max}$, and $T$ are the parameters of the SDE. $c_{\mathrm{niso}}$ and $c_{\mathrm{tango}}$ are the parameters for Niso-DM and Tango-DM, respectively. NN, $n_{0}$, and $\alpha_{\mathrm{ld}}$ are the hyperparameters of the sampling algorithm. $N_{\mathrm{epoch}}$, $B$, $\mathrm{lr}$, and $\mathrm{clip}$ denote the total epochs, the batch size, the learning rate, and the gradient clipping threshold during training. $N_{\mathrm{node}}$ and $N_{\mathrm{layer}}$ represent the number of nodes per layer and the number of hidden layers in the neural networks, respectively. $\mathcal{D}$ indicates the dataset size.
@@
D.1 Riemannian sliced score matching
Diffusion models constrained to a manifold often rely on complex geometric constructs, such as the heat kernel or logarithmic mapping, to compute transition probabilities. These dependencies limit the efficient computation of the denoising score matching loss. In contrast, Riemannian sliced score matching (RSSM [8, 17]) offers broader applicability to general manifolds. This is achieved by leveraging the implicit score matching loss [8], combined with Hutchinson trace estimation [18]. The corresponding training loss is expressed as:
$\ell_{\text{RSSM}}(t,\theta):=\mathbb{E}_{x_{t}}\|s_{\theta}(x_{t},t)\|^{2}+2\mathbb{E}_{z\sim\mathcal{N}(0,I),x_{t}}\left[z^{T}P(x_{t})^{T}\nabla_{x_{t}}s_{\theta}(x_{t},t)P(x_{t})z\right].$ (73)
D.2 The hyperplane in 3D space
Reversal Annealing
[…]
@@ D.2 The hyperplane in 3D space
The target distribution is a mixture of Gaussian distributions with nine modes located on the plane. Specifically, the means of the nine modes are (1,1)(-1,-1), (1,0)(-1,0), (1,1)(-1,1), (0,1)(0,-1), $(0,0)$, $(0,1)$, (1,1)(1,-1), $(1,0)$, and $(1,1)$, and each Gaussian has a standard deviation of $0.3$.
D.3 Mesh data
@@
To create the datasets, we first obtain the clamped eigenfunctions of the Laplacian operator on a mesh that has been upsampled threefold for the original mesh. The target distribution is chosen as an equal-proportion mixture of density functions corresponding to the 0-th, 500-th, and 1000-th eigenpairs. For the distribution on the cow mesh, we additionally discarded the points on the horns and tail.
For $x\notin\mathcal{M}$, the closest point $\pi(x)$ on a triangular mesh surface is determined as follows: First, $x$ is projected onto the planes of all triangular faces in the mesh using the face normals and vertex positions. Barycentric coordinates are then computed to check whether the projected points lie inside the triangles. If a point falls outside a triangle, it is further projected onto the nearest edge to ensure it remains on the triangle’s boundary. Subsequently, the Euclidean distance between the original query point $x$ and all projected points is calculated, and the closest point $\pi(x)$ is selected by identifying the minimum distance.
We assign the normal direction $n(\pi(x))$ of $\pi(x)$ as the normal direction for $x$ ($x\notin\mathcal{M}$). When $\pi(x)$ is located inside a triangle of the mesh, the normal of the triangle is chosen as the normal vector at $\pi(x)$ (i.e. $n(x)=n(\pi(x))$). However, when $\pi(x)$ lies on an edge (or vertex) of the mesh, it is simultaneously associated with two (or three) triangles. In this case, we select the normal vector of the triangle with the smallest index as the normal direction for $\pi(x)$. Here, we leverage the property of the torch.argmin function.
D.4 High-dimensional special orthogonal group
@@
The manifold $\mathrm{SO}(10)$ is defined as $\{Q\in\mathbb{R}^{10\times 10}|QQ^{T}=I_{10},\operatorname{det}(Q)=1\}$, which represents (a connected component of) the zero-level set of the map $\xi:\mathbb{R}^{100}\rightarrow\mathbb{R}^{55}$. The components of $\xi$ correspond to the upper triangular portion of the matrix $QQ^{T}-I_{10}$. The dataset is constructed as a mixture of 5 wrapped normal distributions. Each wrapped normal distribution is the image (under the exponential map) of a normal distribution defined in the tangent space at a pre-selected center. For more details, refer to [23]. In addition, Figure 3 shows that compared with Iso-DM, the samples from Niso-DM have a better match with the ground truth.
Figure 3: Results for $\mathrm{SO}(10)$. Empirical densities of the statistics $\operatorname{tr}(S),\operatorname{tr}(S^{2}),\operatorname{tr}(S^{4})$, and $\operatorname{tr}(S^{5})$ for Iso-DM (red), Niso-DM (purple) and the ground truth (green). Samples from Niso-DM (purple) better match the ground truth (green).
D.5 Alanine dipeptide
@@
To generate the dataset, we follow [23] to obtain the target distribution on the manifold $\mathcal{M}=\{x\in\mathbb{R}^{30}|\phi(x)=-70^{\circ}\}$. We also ensure that the generated distribution is $\mathrm{SE}(3)$-invariant (i.e., invariant under rotations and translations) by incorporating alignment before feeding the data into the neural network.
Appendix E Ablation Study
@@
For Niso-DM, the singularity is governed by the additional noise scale $c_{\mathrm{niso}}$, rather than $\sigma_{\min}$. As a result, our method exhibits greater robustness as $\sigma_{\min}$ decreases. The blue line in Figure 5(a) demonstrates the stability of the error for Niso-DM under varying $\sigma_{\min}$.
Figures 5(b) and 5(c) show the impact of $c_{\mathrm{niso}}$ in Niso-DM and $c_{\mathrm{tango}}$ in Tango-DM on the error. The selection of these hyperparameters also involves a bias-variance trade-off.
@@
In Tables 6-9, we present the numerical results of ablation studies on $\sigma_{\min}$, $c_{\mathrm{niso}}$, and $c_{\mathrm{tango}}$, for the $\mathrm{SO}(10)$ experiment. We report computational errors under different parameter settings to assess the sensitivity of our method to hyperparameter choices. The results demonstrate that reductions in hyperparameter values may not substantially affect the performance of our method.
Table 6: The impact of $\sigma_{\min}$ on the error under the Niso algorithm.
$\sigma_{\min}$ 1e-5 1e-4 5e-4 1e-3 1e-2
[…]
Table 7: The impact of $c_{\mathrm{niso}}$ on the error under the Niso algorithm.
$c_{\mathrm{niso}}$ 1e-3 5e-3 1e-2 5e-2
[…]
Table 8: The impact of $\sigma_{\min}$ on the error under the Tango algorithm.
$\sigma_{\min}$ 1e-5 1e-4 5e-4 1e-3 1e-2
[…]
Table 9: The impact of $c_{\mathrm{tango}}$ on the error under the Tango algorithm.
$c_{\mathrm{tango}}$ 5e-3 1e-2 5e-2 1e-1
[…]
Appendix F Discussion
F.1 Computational cost of Tango-DM
While Tango-DM tends to be slower due to the annealing sampler, the runtime is primarily determined by the number of steps of the inner Langevin dynamics. For example, in the toy experiment R2inR3, the sampling time for Reverse SDE is 1.68 seconds, while for Annealing SDE, the sampling time increases to 5.62 seconds with an inner step of 5 and to 9.40 seconds with an inner step of 10. The first phase of Annealing SDE (see Algorithm 3) requires only 0.82 seconds. Overall, the computational cost of Annealing SDE is approximately 3–6 times higher than that of Reverse SDE. These measurements, conducted on a standard laptop, highlight the relative computational expense.
F.2 Learned manifold case
Unlike previous works on distributions on a known manifold, our method avoids relying on geometric information like geodesics or heat kernels, greatly reducing computational complexity. However, compared to approaches under the manifold hypothesis, our method still requires knowledge of the manifold’s definition, including projection operators. In fields like image and language processing, the manifold structure is often assumed but not explicitly known.
Next, we discuss how to extend our method to learned manifolds, using the AutoEncoder as an example. Specifically, we first train an Encoder $\phi_{\theta_{1}}:\mathcal{M}\to\mathbb{R}^{d}$ and a Decoder $\psi_{\theta_{2}}:\mathbb{R}^{d}\to\mathcal{M}$, which provide a parameterized representation of the manifold. The subspace spanned by $\nabla\psi_{\theta_{2}}(\phi_{\theta_{1}}(x))$ corresponds to the tangent space $T_{x}\mathcal{M}$ at point $x$ on the manifold. Once the basis of the tangent space is obtained, the score function can be further decomposed into its tangential and normal components, which enables the implementation of our proposed methods. Intuitively, the success of this approach hinges on the Autoencoder’s ability to accurately capture the underlying manifold structure. In this work, we leave this approach as a direction for future research.
F.3 Assumptions in Theorem 3.1 and Theorem 4.1
The conclusions in Theorem 3.1 and Theorem 4.1 hold pointwise on the manifold, so in the proof, we only require local boundedness. Specifically, the following assumption is sufficient:
  • For any $x\in\mathcal{M}$, there exists $\delta>0$, such that $\sup_{y\in\mathcal{M}_{x}^{\delta}}\max_{1\leq i,j,j^{\prime}\leq n}\left|\frac{\partial P_{ij}}{\partial y_{j^{\prime}}}(y)\right|<+\infty$, where $\mathcal{M}_{x}^{\delta}=\{\arg\min_{x^{*}\in\mathcal{M}}|\tilde{x}-x^{*}|^{2}\mid|\tilde{x}-x|<\delta\}$.
To make the theorem statement more concise, we adopted a stronger assumption:
  • $\sup_{x\in\mathcal{M}}\max_{1\leq i,j,j^{\prime}\leq n}\left|\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x)\right|<+\infty$.
For compact manifolds, the uniform boundedness assumptions always hold, which implies that the local boundedness assumptions hold. Similarly, the boundedness assumption in Theorem 4.1 can also be relaxed to local boundedness.
Appendix G Impact statement
This paper contributes to the advancement of generative models for data with manifold structures, providing a deeper understanding of the singularity of the score function. Specifically, we identify the scale discrepancies between the tangential and normal components of the score function, which sheds light on key challenges in modeling data on manifolds. We believe that our work bridges the gap between generative models specifically designed for manifolds and those aimed at handling data under manifold assumptions. While this study may have broader societal implications, none require particular emphasis at this stage.